OpenStudy (anonymous):

a sail is in the form of a right triangle that is two timesashigh as it is wide the sail is made from4 square meters of material

OpenStudy (anonymous):

Area = 1/2 base x height 4 square metres = 1/2 base x height but height = 2 x base 4 = 1/2 x base x 2 base 4 = base x base base = square root of 4 = 2 metres Therefore as height = 2 x base, the height = 4 metres